OpenStudy (anonymous):

Express the terms of the following geometric sequence by giving an explicit formula. 6,-30,150,-750, 3,750, ... A. tn=6(-5)(n-1), where n=1,2,3,... B. tn=6(-5)^n-1, where n=1,2,3,... C. tn=6(-5)^n, where n= 1,2,3,... D. tn=-5(6)^n, where n=1,2,3,...

OpenStudy (haseeb96):

C is the correct answer

OpenStudy (anonymous):

here a=6 and r=-5 hence tn=a*r^(n-1)=6*(-5)^(n-1)

OpenStudy (anonymous):

@MarthaHearst

OpenStudy (haseeb96):

@Princer_Jones solve it complete it can never be done that series as given in question statement

OpenStudy (anonymous):

r=2nd term/1st term=3rd term/2nd term=-5 and 1st term is 6 hence use the formula tn=a*r^(n-1)=6*(-5)^(n-1)

OpenStudy (anonymous):

@Haseeb96 you are wrong, according to your solution, check what is the 1st term...

OpenStudy (anonymous):

@Haseeb96 according to you, the 1st term is -30, which is incorrect. do check it